OpenStudy (anonymous):

what are the x intercepts of 4(x+1)^2-2

OpenStudy (anonymous):

unless its equal to y there are no x-intercepts

OpenStudy (anonymous):

agree with ^^^. do you mean \(\Large y=4(x+1)^2-2 \) ???

OpenStudy (anonymous):

it says exactly is find the x intercepts of f(x)=4(x+1)^2-2

OpenStudy (cwrw238):

when f(x) = 0 4(x+1)^2-2 = 0 solve for x

OpenStudy (anonymous):

ok.... so y=f(x)... to find the x-intercepts, set y=0, or set your quadratic equal to zero and solve for x...